Calculate the angular displacement of an object rotating from pi/2 to 4pi/3
1
Expert's answer
2020-04-13T10:09:52-0400

The angular displacement of an object:


"\Delta \theta=\theta_2-\theta_1"

"\Delta \theta=\frac{4\pi}{3}-\frac{\pi}{2}=\frac{5\pi}{6}\ rad=150\degree"